1−

2[Campodedeslocamentosdabaseposterior.V istadefrente]Campodedeslocamentosdocomponentebaseposteriorsujeitoaumacargade.V istadefrente. 1−

2_{1}[Deformadadabaseposterior]Deformadadabaseposterioraumaescalade100vezesovalorrealdodeslocamento.

ten_{b}ase_{t}ras[Campodetensesdabaseposterior]Campodetensesdocomponentebaseposteriorsujeitoaumacargade. 2−

2[Campodedeslocamentosdabaseanterior.V istadefrente]Campodedeslocamentosdocomponentebaseanteriorsujeitaaumacargade.V istadefrente.×

2− 2_{1}[Deformadadabaseanterior]Deformadadabaseanterioraumaescalade100vezesovalorrealdodeslocamento. ten_{b}ase_{f}ren[Campodetensesdabaseanterior]Campodetensesdocomponentebaseanteriorsujeitoaumacargade.

5_{5}_{1}[Campodedeslocamentosedeformadadocomponentedobrao]Campodedeslocamentosdocomponentedobraosujeitoaumacargadeerespectivadeformadaaumaescalade1000vezesovalorrealdodeslocamento. ten_{b}raco_{1}[Campodetensesdocomponentedobrao]Campodetensesdocomponentedobraosujeitoaumacargade.

6_{6}_{1}[Campodedeslocamentosedeformadadocomponentedobrao]Campodedeslocamentosdocomponentedobraosujeitoaumacargadeerespectivadeformadaaumaescalade1000vezesovalorrealdodeslocamento. 7_{7}_{2}[Definiodascondiesimpostasecampodedeslocamentosdocomponentedobrao]Definiodascondiesimpostasecampodedeslocamentosdocomponentedobraosujeitoaumacargade. 7_{1}[Deformadadocomponentedobrao]Deformadadocomponentedobraoescalarealdodeslocamento. 8_{8}_{2}[Definiodascondiesimpostasecampodedeslocamentosdocomponentedobrao]Definiodascondiesimpostasecampodedeslocamentosdocomponentedobraosujeitoaumacargade. 8_{1}[Deformadadocomponentedobrao]Deformadadocomponentedobraoescalarealdodeslocamento.

9[Campodedeslocamentosdocomponentedobrao]Campodedeslocamentosdocomponentedobraosujeitoaumacargade. 9_{1}[Deformadadocomponentedobrao]Deformadadocomponentedobraoaumaescalade10vezesovalorrealdodeslocamento.

10[Campodedeslocamentosdocomponentedobrao]Campodedeslocamentosdocomponentedobraosujeitoaumacargade. 10_{1}[Deformadadocomponentedobrao]Deformadadocomponentedobraoaumaescalade10vezesovalorrealdodeslocamento.

11_{1}1_{1}[Campodedeslocamentosedeformadadocomponentedobrao]Campodedeslocamentosdocomponentedobraosujeitoaumacargadeerespectivadeformadaaumaescalade10vezesovalorrealdodeslocamento.